OpenStudy (anonymous):

Reduce the fraction 20/100

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

\[ \frac{20}{100}=\frac{2*2*5}{2*2*5*5}=\frac15 \]

OpenStudy (anonymous):

or divde both by twenty because 100 is divisible by 20 :) and you get the same thing. 20 goes into 20 1 time 100 goes into 20 5 times. therefore it equals 1/5 :)